OpenStudy (anonymous):

Explain, in complete sentences, how you would completely factor 15x2 – 25x – 60 and check your factors for accuracy.

OpenStudy (anonymous):

15x^2-25x-60=0 5(3x^2-5x-12)=0 3x^2-5x-12=0 3x^2-9x+4x-12=0 3x(x-3)+4(x-3)=0 x=3,-4/3